While putting, Kevin made 16 out of his last 21 putts. His friend Mike made 9 out of his last 14 putts. What is the probability that they both make their next putt? Leave answer as a simplified fraction.

 




1
Expert's answer
2022-02-01T13:02:09-0500

Solution:

Probability that Kevin makes putts",P(K)=\dfrac{16}{21}"

Probability that Mike makes putts",P(M)=\dfrac{9}{14}"

Probability that both make putts",P(K\cap M)=\dfrac{16}{21}\times\dfrac{9}{14}=\dfrac{24}{49}"